f(x,y)=e^(2^y2-x^2+4y) 1.what is fxx fxy and fyy? 2. use the method of Lagrange multiplier to find local max and min of f(x,y)=x^2-y sbuject to constraint g(x,y)=x^2+y^2-1=0.
Answer #1

Date: x²+y2=1 7 ? + +(%)? => x ² +/4 = 1 So 7 m² = 1-44= 3/4 = x2 = / 7 = + / / the critical point ther f( + 2 1 1,-/h) = 22femy), la fex +49) x² + 4y)= 22 % pr++48) for = {2+2x+43) 24 2 Pant (2nd com - e 4+48) x (-2) -2.0. from 2y -ala: 62% +40) *24 (2 lena. 2+4) 28 (22 x ² +40) fy s diena.e fay 27 (29p++49) (22x2449) (-2x) + 4.e +46 (2) arthy) 24 = 2lna.ee - 48.lmp.ey.